OpenStudy (anonymous):

hard question x^2-16over x^2+5x+6 divided by x^2+4xover x^2+6x+9

OpenStudy (anonymous):

show steps if possible

OpenStudy (lalaly):

\[\frac{\frac{(x-4) \cancel{(x+4)}}{\cancel{(x+3)}(x+2)}}{\frac{x \cancel{(x+4)}}{\cancel{(x+3)}(x+3)}}\]\[=\frac{(x+3)(x-4)}{x(x+2)}\]